Big Apple Cabins has net income of $725,000. Throughout the year, the company had 150,000 shares of common stock outstanding. Also, the company has 25,000 shares of preferred stock that pay a dividend of $5.00 per share that is convertible into 5 shares of common stock for each share of preferred. The preferred stock is considered to be dilutive. The tax rate for Big Apple Cabins is 40%. What are the diluted earnings per share for Big Apple Cabins?

Select one: A. $2.64 B. $4.00 C. $2.25 D. $2.50 E. $2.00
Answer #1

Basic earning per share = 725000-25000*5/150000 = 4.00

Diluted earning per share = Adjusted net income/Adjusted share = 725000/125000+150000 = 2.64

So answer is a) $2.64